OpenStudy (sighn0more):

Can someone help me work through this problem? Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8

OpenStudy (anonymous):

Get them into slope-intercept form first

OpenStudy (sighn0more):

Okay, one second......

OpenStudy (sighn0more):

4y-x=-8 in slope-intercept form is y=x/4-2. I'm almost done with the second one.....

OpenStudy (anonymous):

it would actually be y=1/4x-2 because you have to pick up the value of x which is one

OpenStudy (loser66):

The problem ask you to draw them first, now do it Can you?

OpenStudy (loser66):

Don't for get!! GRAPH the equations to SOLVE them. Not solve in algebraic way.

OpenStudy (loser66):

Let me know if you want to know how to "GRAPH to SOLVE"

OpenStudy (sighn0more):

I think that I get it now! :) I converted the second equation to slope-intercept form and I found that both of the equations were the same. So I think that this line has infinitely many solutions.

OpenStudy (anonymous):

You have it :)

OpenStudy (sighn0more):

Thank you both!

OpenStudy (anonymous):

You're welcome